OpenStudy (loser66):

The last one for tonight If A, B, C, D are linear transformations (all on the same vector space)and if both (A+B) and (A-B) are invertible, then there exist linear transformations, X and Y such that AX + BY = C and BX +AY = D Please, I don't get what I am supposed to do

OpenStudy (loser66):

Do they ask me to prove there exists X and Y like that?

OpenStudy (ikram002p):

yess m but u have the key, start from invertible,

OpenStudy (loser66):

hehehe. not too far .

OpenStudy (ikram002p):

so got it or should we do it here ?

OpenStudy (loser66):

Please, give me some steps. I am so tired for those stuff

OpenStudy (anonymous):

start by adding & subtracting the 2 equations

OpenStudy (loser66):

@eashy We can't use those equations, they are conclusion we have to prove, right?

OpenStudy (ikram002p):

invertible gives u (A+B)(A-B)=I G cong (G)=I G=cong ^-1 (G) let me remember what we conclude from that

OpenStudy (anonymous):

we can assume the equations are true for now, and then use the equations to find explicit solutions for X and Y

OpenStudy (anonymous):

thereby proving the equations

OpenStudy (loser66):

@ikram002p (A+B) and (A-B) invertible do not mean (A+B) is inverse of (A-B)

OpenStudy (anonymous):

1) add the equations: (A+B)(X+Y)=C+D 2) subtract the equations: (A-B)(X-Y)=C-D 3) solve for X,Y

OpenStudy (loser66):

@eashy oh yes, I got what you mean, hehehe

OpenStudy (ikram002p):

hought it mean (A+B) (A-B)=(A-B)(A+B)=I right ??!

OpenStudy (loser66):

@ikram002p I don't think so, consider (A+B) = T, so \(T^{-1} \) totally different from inverse of (A-B)

OpenStudy (ikram002p):

hmm ok ! my bad lol

OpenStudy (loser66):

No problem, You help me a lot. hihihi

OpenStudy (loser66):

@eashy Thank you so much. It's 2:00 am, now I have to go to bed.

OpenStudy (ikram002p):

lol its 6:54 here

OpenStudy (loser66):

where are you from?